OpenStudy (stephend95):

What is the approximate solution of this system of equations? y = x2 + 3x + 3 y = x2 + x + 2 A. (-0.5, 0.2) B. (-0.5, 1.7) C. (-0.1, 1.2) D. (0.2, 1.5)

OpenStudy (stephend95):

I honestly have no idea what this question is even asking. Anyone who can help me one this i really would appreciate it, thanks and godbless

OpenStudy (1davey29):

A solution to a system of equations is a point in which they intersect. Because they are both y=, x^2+3x+3=x^2+x+2. Solve for x, subtract x^2 and x from both sides, 2x+3=2, subtract 3 from both sides, 2x=-1, divide by 2, x=-(1/2). To find your y value, plug x into either equation. For simplicity, I will use the bottom one. (-1/2)^2+(-1/2)+2=(1/4)-(1/2)+2=1 3/4. Convert both to decimal, .5 and 1.75, or 1.7, to get (0.5, 1.7), or B

OpenStudy (stephend95):

Thank you brother. I really appreciate man, have a good day.